# Variational Quantum Eigensolvers with Quantum Gaussian Filters for solving ground-state problems in quantum many-body systems

### 24 Jan 2024 v1

@@ I Introduction
The paper is structured as follows: Section II details the principles and procedures of our quantum algorithm. The algorithm’s efficacy in solving the ground-state problem is demonstrated numerically in Section III. Finally, Section IV concludes the paper and discusses future implications.
II Quantum Algorithm
@@
VQE Implementation for Step Evolution
@@
$C(\vec{\theta}^{(t+1)})=-|\braket{\psi(\vec{\theta}^{(t+1)})}{\psi(\vec{\theta}^{(t)})}-\Delta\tau\braket{\psi(\vec{\theta}^{(t+1)})}{\hat{H}^{2}}{\psi(\vec{\theta}^{(t)})}|.$ | . (2)
@@
Incorporating McLachlan’s Variational Principle in VQE
@@
Procedure of the Quantum Algorithm
@@
III Numerical Demonstration
In this section, we provide a numerical demonstration of our algorithm designed for solving the ground-state in quantum many-body models. Our focus is primarily on the Transverse Field Ising Model. The simulations are conducted using the open-sourced quantum simulation package QuTiP [47]. The Hamiltonian of the Transverse Field Ising Model [48] is represented as:
@@ III Numerical Demonstration
where $J$ is the interaction strength of nearby sites, $g$ is the scale of the external transverse field, and $\hat{\sigma}^{z}_{N+1}=\hat{\sigma}^{z}_{1}$ for periodic boundary condition. Two sets of parameters are investigated: $J=0.5,g=1$, which leads to a dominant $g$ term indicative of the paramagnetic phase, and $J=1,g=0.5$, where the $J$ term prevails, signifying the ferromagnetic phase.
@@
We select a system size of $N=4$, and for both two cases, we apply an energy equaling to $8.5$ shift the ground-state to around $4.228$. The initial state for each phase is prepared with specific considerations. For the ferromagnetic phase, characterized by spin alignment, each qubit is initialized in the ground state of $\hat{\sigma}^{x}$ by applying a Pauli-Z gate after a Hadamard gate to |$\ket{0}$, i.e. |$\ket{\psi_{i}}=\otimes_{n=1}^{N}\hat{\sigma}_{n}^{z}H\ket{0}$. In the paramagnetic phase, where spins align more with the external field, we create a GHZ state using Hadamard and CNOT gates, followed by a phase-flip gate on each qubit. A $p=4$ layer QAOA ansatz [49] is employed. The parameterized state is described as |$\ket{\psi(\vec{\theta})}=[\prod_{p=0}^{3}e^{-i\theta_{2p}\hat{h}_{x}}e^{-i\theta_{2p+1}\hat{h}_{zz}}]\ket{\psi_{i}}$, where $\hat{h}_{zz}$ and $\hat{h}_{x}$ correspond to the first and second terms of Eq. (5), respectively, and |$\ket{\psi_{i}}$ is the prepared initial state.
@@
The initial parameters $\vec{\theta^{(0)}}$ are generated randomly, with 50 distinct sets for each simulation. We evolve the quantum Gaussian filter for $\tau=0.2$ with a step size of $\Delta\tau=0.005$, resulting in a total of 40 evolution steps. At each step, we minimize the cost function Eq. (2) using gradient descent, limiting iterations to a maximum of 10 per step. If there is no improvement in cost, the training process is interrupted, and we proceed to the next evolution step.
@@
Fig. 1 displays the results of our algorithm. It shows the mean and standard deviation of the estimated ground-state energy at each evolution step. Additionally, it presents the fidelity between the algorithmically prepared state and the exact ground state. These results affirm that our method effectively evolves different initial states towards the target ground state.
Figure 1: Numerical results for the Transverse Field Ising Model with $N=4$. The yellow line represents the mean value of the estimated ground-state energy. The blue error bars indicate the standard deviation of the estimated energy. The red dashed line marks the exact ground-state energy, providing a reference for comparison. The purple line depicts the mean fidelity between the prepared state and the exact ground state. (a) illustrates the model in the ferromagnetic phase, while (b) shows the model in the paramagnetic phase.
@@
In our investigation, we also analyze the impact of varying shift-energy values. According to Eq. (1), the effective weighting of various eigenstates, including the ground state, is influenced by the energy scale. An increase in shift-energy generally reduces the weight of additional eigenstates while accentuating the relative importance of the ground state. This adjustment can potentially enhance the performance of the algorithm. However, a higher shift-energy also necessitates approximating the QGF filter over a broader energy range, requiring a smaller step size $\Delta\tau$. In this demonstration, we employ the shift-energy values of $4.5$, $5.5$, and $6.5$, aiming to adjust the ground-state energy around $0.228$, $1.228$, and $2.228$, respectively. We set $\Delta\tau=0.002$ and evolve the QGF to $\tau=0.15$. For parameter updates, we utilize the McLachlan method. Fig. 2 presents our numerical results, showcasing the estimated energy error as a function of evolution $\tau$. These findings indicate that a higher shift-energy leads to a faster convergence rate, aligning with our theoretical expectations.
@@ IV Conclusion
Numerical demonstrations on the Transverse Field Ising Model and the Heisenberg Model substantiate the algorithm’s potential. Our simulations, including scenarios with depolarizing noise, exhibit promising results. The algorithm consistently produces states with high fidelity to the exact ground states, demonstrating its robustness and adaptability to various quantum many-body models and noisy environments. Compared to trivial VQE method, our approach has a faster converge rate and a higher expected fidelity. Notably, it performs exceptionally well when initiated from a well-prepared state, suggesting broader applicability to other scenarios with similarly prepared initial states, as discussed in Ref. [50, 51, 52, 1].
@@
Acknowledgements.
We thank Dan-Bo Zhang, Yan Zhu, and Yu-Cheng Chen for valuable discussions.

### 15 Mar 2024 v2

@@ 1 Introduction
The paper is structured as follows: Section 2 details the principles and procedures of our quantum algorithm. The algorithm’s efficacy in solving the ground-state problem is demonstrated numerically in Section 3. Finally, Section 4 concludes the paper and discusses future implications.
2 Quantum Algorithm
@@ 2.1 QGF algorithms
2.2 VQE Implementation for Step Evolution
@@
$\begin{split}C(\vec{\theta}^{(t+1)})=&-|\braket{\psi(\vec{\theta}^{(t+1)})}{\psi(\vec{\theta}^{(t)})}\\ &-\Delta\tau\braket{\psi(\vec{\theta}^{(t+1)})}{\hat{H}^{2}}{\psi(\vec{\theta}^{(t)})}|.\end{split}$ (2)
@@
2.3 Incorporating McLachlan’s Variational Principle in VQE
@@
2.4 Procedure of the Quantum Algorithm
@@
3 Numerical Demonstration
Figure 1: Numerical results for the Transverse Field Ising Model. The yellow line represents the mean value of the estimated ground-state energy. The blue error bars indicate the standard deviation of the estimated energy. The red dashed line marks the exact ground-state energy, providing a reference for comparison. The purple line depicts the mean fidelity between the prepared state and the exact ground state. (a) illustrates the model with $N=4$ qubits in the ferromagnetic phase, (b) shows the model with $N=4$ qubits in the paramagnetic phase, (c) illustrates the model with $N=6$ qubits in the ferromagnetic phase, while and (d) shows the model with $N=6$ qubits in the paramagnetic phase.
In this section, we provide a numerical demonstration of our algorithm designed for solving the ground state in quantum many-body models. Our focus is primarily on the Transverse Field Ising Model. The simulations are conducted using the open-sourced quantum simulation package QuTiP [47]. The Hamiltonian of the Transverse Field Ising Model [48] is represented as:
@@ 3 Numerical Demonstration
where $J$ is the interaction strength of nearby sites, $g$ is the scale of the external transverse field, and $\hat{\sigma}^{z}_{N+1}=\hat{\sigma}^{z}_{1}$ is for periodic boundary condition. Two sets of parameters are investigated: $J=0.5,g=1$, which leads to a dominant $g$ term indicative of the paramagnetic phase, and $J=1,g=0.5$, where the $J$ term prevails, signifying the ferromagnetic phase.
@@
We select system sizes of $N=4$ and $N=6$. For both cases, we apply an energy shift of $8.5$ to move the ground-state energy to approximately $4.228$ for the $4$-qubit case and $2.115$ for the $6$-qubit case. The initial state for each phase is prepared with specific considerations. For the ferromagnetic phase, characterized by spin alignment, each qubit is initialized in the ground state of $\hat{\sigma}^{x}$ by applying a Pauli-Z gate after a Hadamard gate to |$\ket{0}$, i.e. |$\ket{\psi_{i}}=\otimes_{n=1}^{N}\hat{\sigma}_{n}^{z}H\ket{0}$. In the paramagnetic phase, where spins align more with the external field, we create a GHZ state using Hadamard and CNOT gates, followed by a phase-flip gate on each qubit. A $p=4$ layer QAOA ansatz [49] is employed. The parameterized state is described as |$\ket{\psi(\vec{\theta})}=[\prod_{p=0}^{3}e^{-i\theta_{2p}\hat{h}_{x}}e^{-i\theta_{2p+1}\hat{h}_{zz}}]\ket{\psi_{i}}$, where $\hat{h}_{zz}$ and $\hat{h}_{x}$ correspond to the first and second terms of Eq. (5), respectively, and |$\ket{\psi_{i}}$ is the prepared initial state.
@@
The initial parameters $\vec{\theta^{(0)}}$ are generated randomly, with $50$ distinct sets for each simulation. We evolve the quantum Gaussian filter for $\tau=0.15$ with a step size of $\Delta\tau=0.005$, resulting in a total of $30$ evolution steps. At each step, we minimize the cost function Eq. (2) using gradient descent, limiting iterations to a maximum of $10$ per step. If there is no improvement in cost, the training process is interrupted, and we proceed to the next evolution step.
@@
Fig. 1 displays the results of our algorithm. It shows the mean and standard deviation of the estimated ground-state energy at each evolution step. Additionally, it presents the fidelity between the algorithmically prepared state and the exact ground state. These results affirm that our method effectively evolves different initial states towards the target ground state, and can handle the scale effect as well.
@@
In our investigation, we also analyze the impact of varying shift-energy values. According to Eq. (1), the effective weighting of various eigenstates, including the ground state, is influenced by the energy scale. An increase in shift-energy generally reduces the weight of additional eigenstates while accentuating the relative importance of the ground state. This adjustment can potentially enhance the performance of the algorithm. However, a higher shift-energy also necessitates approximating the QGF filter over a broader energy range, requiring a smaller step size $\Delta\tau$. In this demonstration, we employ the $N=4$ qubit model with shift-energy values of $4.5$, $5.5$, and $6.5$, aiming to adjust the ground-state energy around $0.228$, $1.228$, and $2.228$, respectively. We set $\Delta\tau=0.002$ and evolve the QGF to $\tau=0.15$. For parameter updates, we utilize the McLachlan method. Fig. 2 presents our numerical results, showcasing the estimated energy error as a function of evolution $\tau$. These findings indicate that a higher shift-energy leads to a faster convergence rate, aligning with our theoretical expectations.
@@ 4 Conclusion
Numerical demonstrations on the Transverse Field Ising Model substantiate the algorithm’s potential. Our simulations, including scenarios with depolarizing noise, exhibit promising results. The algorithm consistently produces states with high fidelity to the exact ground states, demonstrating its robustness and adaptability to various quantum many-body models and noisy environments. Compared to trivial VQE method, our approach has a faster converge rate and a higher expected fidelity. Notably, it performs exceptionally well when initiated from a well-prepared state, suggesting broader applicability to other scenarios with similarly prepared initial states, as discussed in Ref. [50, 51, 52, 1].
@@
Acknowledgements
We thank Dan-Bo Zhang, Yan Zhu, and Yu-Cheng Chen for valuable discussions. This work was supported by the Guangdong-Hong Kong Joint Laboratory of Quantum Matter, the NSFC/RGC JRS grant (RGC Grant No. N_HKU774/21, NSFC Grant No. 12161160315), the CRF (Grant No. C6009-20G) and GRF (Grant No. 17310622) of Hong Kong.